OpenStudy (anonymous):

Make the following conversions: A. 42.7 L to milliliters B. 261 nm to milliliters C. 0.065 km to decimeters Please explain your answers so I can understand how you got them..

OpenStudy (anonymous):

a) 42.7L / 1L * 1000mL = 42700mL There's 1000 mL in 1 L b)261nm/10^9nm*1m=2.61*10^-7m There are 10^9 nm in 1 m c)0.065 km/1km*1000m*10dm/1m =650dm There is 10 dm in 1 m, and 1000 m in 1km

OpenStudy (anonymous):

@ajmayberry thank you so much!